Dear Tutor,
These are the problems I'm having trouble with...
At the same time of day a man who is 76inches tall casts a 57inch shadow and his son casts a 24inch shadow. What is the height of the son? (I've tried over and over and can't get it right! Please help ASAP thank you very much!)
Also.....
Given triangle DEF, AB=5, BC=10, DE=8, and DF=2. find the Ratio of the lengths of the corresponding sides of triangle ABC to triangle DEF. ( this one just doesn't make sense to me!)
These are my troubled problems...please help me ASAP it's due tomorrow:0 sorry to dump this on you! Thank you so so soooooo very much!
-from Hopeless

Answer by sudhanshu_kmr(1152)


The ratio of height and length of the shadow will always equal for men and his
son.

i.e height of son / length of shadow = height of man / length of his shadow
=> height of son/24 = 76/57
=> height of son = 76/57 * 24 = 32
so, height of son = 32 inch

In second question, something is missing. Question is not clear.

You are welcome to contact me if any doubt...